OpenStudy (anonymous):

solve for x. 14+ x/5 =17.25?

OpenStudy (mertsj):

First multiply every term by 5 to clear the fractions. Do that and tell me what you get.

OpenStudy (anonymous):

mutiply 5 on both sides

OpenStudy (anonymous):

the final answer is x=16.26

OpenStudy (anonymous):

its like 5*14+x/5*5=17.25*5

OpenStudy (anonymous):

then solve for x

OpenStudy (anonymous):

subtract 14 from 14 it crosses off

OpenStudy (anonymous):

then subtract 17. 25 - 14 which is 3.25

OpenStudy (anonymous):

then you do the opposite multiply 3.25 * 5 = 16.25

OpenStudy (anonymous):

16.25 is your answer

OpenStudy (anonymous):

x=16.25*

OpenStudy (anonymous):

thats how you really do it

OpenStudy (anonymous):

check if you want by putting it in the eqaution

OpenStudy (anonymous):

good luch with your studies